OpenStudy (anonymous):

Megan and Sam together take 12 minutes to wash a car. Washing the car by herself Megan took 21 minutes. How long would it take Sam to wash the car? This is how I started: 1/12 = 1/21 + 1/X 1/12 - 1/21 = 1/X Now what should I do ?? Cross Multiply?

OpenStudy (anonymous):

hmm, this is a tricky question, i'm gonna have to think out side the box, give me a moment.

OpenStudy (anonymous):

sure take your time :)

OpenStudy (anonymous):

wow this actually stumped me, but i would assume that somewhere there is going to be a half in there.

OpenStudy (anonymous):

i'd guess it took her 27 minutes, but that is a total wide guess.

OpenStudy (anonymous):

@luigi help out budy

OpenStudy (anonymous):

@Luigi0210 my bad

OpenStudy (anonymous):

Thanks for trying

OpenStudy (zeeporo):

i would approach from the speed of the persons. Megan x, Sam y.

OpenStudy (zeeporo):

if they make together speed is x+y. so doing the work 1/(x+y) = 12

OpenStudy (zeeporo):

make sense?

OpenStudy (anonymous):

no

OpenStudy (zeeporo):

let me sort out...

OpenStudy (zeeporo):

I got it, do you want to think it over, or just the result?

OpenStudy (anonymous):

i want to know how you got the result my test is tomorrow

OpenStudy (zeeporo):

OK, as I wrote approach it from the speed of the persons. Just this is not KM/hour but 1 carwash / minute. Megan speed is x. And we know that it takes her 21 minutes, this means 1/x = 21.

OpenStudy (zeeporo):

Sam speed is y and we are looking for 1/y which will give us the time (in minutes) he washed the car.

OpenStudy (zeeporo):

If we assume that their speed adds together as it is we know when they work together they washed in 12 minutes 1/(x+y) = 12.

OpenStudy (zeeporo):

Is this logic so far?

OpenStudy (anonymous):

yes

OpenStudy (zeeporo):

OK, now you only have to solve the equations. What is your result?

hero (hero):

If Megan can wash the car in M minutes while working alone Sam can wash the car in S minutes while working alone Then Together, they can wash the car in t minutes while working alone As represented by the following equation: \[\frac{MS}{M + S} = t\] Now in this case M = x (since it is not given) S = 21 t = 12 (which means they can do the job together in 12 minutes) So therefore, plugging the given info in to the equation, we have: \[\frac{21x}{21 + x} = 12\] Solve for x

hero (hero):

Solving for \(x\) will give how long it takes for Megan to wash the car while working alone.

hero (hero):

\(21x = 12(21 + x)\) \(21x = 252 + 12x\) \(21x - 12x = 252\) \((21 - 12)x = 252\) \((20 + 1 - 10 - 2)x = 252\) \((20 - 10 + 1 - 2)x = 252\) \((10 - 1)x = 252\) \(9x = 252\) \(x = 252/9\)

OpenStudy (zeeporo):

Do you have the result? Same as Hero's.

OpenStudy (anonymous):

yup thanks guys

OpenStudy (wolf1728):

the formula for total time is: T = M * S / (M+S) 12 = 21 * S / (21 + S) 12 * (21 + S) = (21 * S) 252 + 12S = 21S 252 = 9S S = 252 / 9 minutes S = 28 minutes

OpenStudy (anonymous):

Thanks for all the help !!

OpenStudy (zeeporo):

good luck on the test

OpenStudy (wolf1728):

u r welcome

hero (hero):

I seem to have mixed up the variables.